OpenStudy (anonymous):

Prove that the diagonals of the parallelogram formed by 4 straight line are at right angles to each other. \[L1=\sqrt{3}x+y=0\] \[L2=x+\sqrt{3}y=0\] \[L3=\sqrt{3}x+y=1\] \[L4=x+\sqrt{3}y=1\]

OpenStudy (anonymous):

i know one way by finding all the corner coordinates.

OpenStudy (anonymous):

but we have to solve it using family of lines.

OpenStudy (anonymous):

@mathslover @ganeshie8 @hartnn @Hero

mathslover (mathslover):

can we say that : \(\large{\sqrt{3}{x}+y=0=\sqrt{3}{x}+y=1}\) ???

OpenStudy (anonymous):

no

mathslover (mathslover):

is \(\large{\sqrt{3}x+y=\sqrt{3}{x}+y}\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

can u find angle between lines ? do u need to use pair of lines equation ? ax^2+2hxy+by^2 ??

mathslover (mathslover):

given : \ L1 = \(\sqrt{3}{x+y=0}\) L3 = \(\sqrt{3}{x}+y=1\) How is it possible?

OpenStudy (anonymous):

parallels

hartnn (hartnn):

those are equatiion of lines

mathslover (mathslover):

0 = 1?

OpenStudy (anonymous):

yes

mathslover (mathslover):

OK gotcha

mathslover (mathslover):

Sorry :)

ganeshie8 (ganeshie8):

(0,0) is one corner as both adjacent lines pass thru origin (y = mx) and (0,1) is another corner as other set of adjacent lines has same y-intercept = 1

ganeshie8 (ganeshie8):

so the slope of one diagonal is undefined

ganeshie8 (ganeshie8):

we need to prove slope of other diagonal = 0

OpenStudy (anonymous):

is there any other method which does not involves finding the corners??

ganeshie8 (ganeshie8):

let me think.. .

OpenStudy (experimentx):

a rectangle is a parallelogram but doesn't have perpendicular diagonals. could you revise your question.

OpenStudy (anonymous):

a square and rhombus are also parallelogram

OpenStudy (experimentx):

you can say for parallelogram them, parallelogram means rectangle and sheared rectangle too .... and square is a rhombus. perhaps you should have said for rhombus.

OpenStudy (experimentx):

also your approach works ... if you are taking those 4 lines as 4 sides of rhombus then it should work fine. also you could try taking 2 diagonals directly as diagonals will be bisecting two sides from on the point they meet.

OpenStudy (anonymous):

i just want to know , how can we find the equations of diagonals from eq. of 4 sides , without finding the corner points??

OpenStudy (experimentx):